if secx=4 what does cos2x=
Answer by jim_thompson5910(35256)

sec(x) = 4

1/cos(x) = 4

1 = 4*cos(x)

1/4 = cos(x)

cos(x) = 1/4

-------------------------------------------------------

cos(2x)

2*(cos(x))^2 - 1

2*(1/4)^2 - 1

2*(1/16) - 1

1/8 - 1

-7/8


So if sec(x) = 4, then cos(2x) = -7/8